OpenStudy (anonymous):

Let f(x) = 3x + 5 and g(x) = x - 4. Find f o g and its domain Answer choices 3x + 1; all real numbers 3x + 1; all real numbers except x = 4 3x - 7 ; all real numbers 3x - 17; all real numbers except x = 4

OpenStudy (accessdenied):

(f o g)(x) = f( g(x) ) = 3(x - 4) + 5 = 3x - 12 + 5 = 3x - 7 There are no restrictions on the domain with this function, so all real numbers are applicable.